# Reshaping Neural Representation via Associative, Presynaptic Short-Term Plasticity

### v1

@@ 4 Discussion
We have demonstrated that extending Fisher information optimization to synapses with Tsodyks–Markram dynamics yields learning rules that preferentially strengthen anti-causal connections, where presynaptic neurons lag behind postsynaptic neurons. This temporal asymmetry emerges from the interaction between associative learning and short-term depression, creating neural representations that emphasize stimulus offsets through ramp-up dynamics. The degree of temporal asymmetry depends critically on constraints governing vesicle release probability. As we discuss below, this combination of anti-causal connectivity and release-probability constraints provides a plausible computational account of state-dependent differences in hippocampal replay directionality observed during wakefulness versus sleep.
@@ 4.1 Pre- versus postsynaptic plasticity: Computational division of labor
The segregation of synaptic plasticity mechanisms between pre- and postsynaptic sites has long served as a guiding principle for understanding synaptic computation. Classical long-term potentiation is mediated primarily by postsynaptic mechanisms, notably changes in AMPA receptor trafficking and expression [malenka2004LTP]. By contrast, short-term plasticity has typically been modeled as a predominantly presynaptic phenomenon governed by vesicle dynamics and calcium-dependent modulation of release probability [zucker2002Shortterm, regehr2012ShortTerm]. This division of labor has been interpreted as separating associative learning (postsynaptic, detecting pre-post correlations) from adaptive filtering (presynaptic, responding only to presynaptic history).
Recent theoretical work has argued that this pre–post division reflects a deeper computational separation rather than a purely biophysical contingency. Postsynaptic mechanisms, operating on longer timescales, are well suited for stable memory storage and the formation of neural assemblies that encode persistent representations [fusi2005Cascade, zenke2015Diverse]. Presynaptic mechanisms, with their rapid dynamics, can implement adaptive filters that extract temporal features from spike trains, enhance signal-to-noise ratios, and maintain homeostatic balance in network activity [abbott2004Synaptic, fortune2001Shortterm].
The discovery of associative short-term plasticity by Ucar et al. [ucar2021Mechanical] challenges this traditional dichotomy. Their experiments demonstrate that presynaptic release probability can be modulated via postsynaptic spine enlargement, in a manner that requires coincident pre- and postsynaptic activity. This mechanism therefore combines the associative detection capabilities conventionally attributed to postsynaptic plasticity with the rapid, history-dependent filtering characteristic of presynaptic dynamics. Our theoretical framework indicates that such a hybrid rule is not simply “LTP made faster”, but gives rise to distinct forms of temporal selectivity.
@@ 4.2 Reverse replay and state-dependent plasticity
This mechanism aligns with and extends the model proposed by [haga2018Recurrent], which demonstrated that symmetric STDP combined with short-term depression can generate backward connectivity supporting reverse replay. In our framework, such connectivity patterns arise from an information-theoretic optimization principle: a backward bias is not an incidental artifact of a particular plasticity rule, but an expected outcome when information transmission is optimized under appropriate constraints.
@@
We interpret this “release-probability budget” as an effective parameter that can be influenced by physiological mechanisms such as neuromodulation. Neuromodulators vary strongly across behavioral states and sleep stages and modulate presynaptic properties, including short-term dynamics [hasselmo2006Role, sara2009Locus]. Recent work further underscores that presynaptic resources are regulated at the behavioral-state level: [wu2025Presynaptic] showed across multiple circuits in Drosophila and mice that presynaptic release probability decreases with prolonged wakefulness and recovers after sleep, and that bidirectional manipulation of release probability bidirectionally alters sleep need. Complementary studies in Drosophila demonstrate that presynaptic active zone composition, which closely tracks release probability, is remodeled as a function of sleep pressure [huang2020Presynaptic]. Although these findings do not address hippocampal replay directly, they establish that presynaptic resource constraints can indeed be tuned in a state-dependent manner, consistent with the key premise of our model.
@@
This picture dovetails with empirical observations that reverse replay is especially prominent during quiet wakefulness following spatial experience and reward delivery, whereas forward replay is more prevalent during sleep sharp-wave ripples [foster2006Reverse, diba2007Forward, ambrose2016Reverse, buzsaki2015Hippocampal, olafsdottir2018Role]. In our framework, these regimes correspond to different effective presynaptic budgets acting on top of a fixed forward backbone: more generous budgets promote the emergence of reverse replay that may facilitate rapid behavioral adaptation and reward credit assignment [foster2006Reverse], whereas more stringent budgets yield connectivity that expresses predominantly forward replay and may better support stabilization and generalization of memories during sleep [gais2004Declarative, rasch2013Sleeps, buzsaki2015Hippocampal, olafsdottir2018Role]. This proposal links state-dependent neuromodulation of presynaptic function to a quantitative, information-theoretic control knob for the balance between forward and reverse replay, and thereby frames reverse sequences not as an oddity of hippocampal dynamics, but as a predicted feature of efficient temporal coding under varying resource constraints.
@@ 4.3 Frequency-Dependent Phase Selectivity in Pre- and Postsynaptic Channels
Our linear-response analysis of $C^{Z}(t)$ (Figure 2; Appendix D.2) confirms that onset selectivity is robust, but the phase of maximal plasticity shifts with input frequency and the neuron–synapse operating point. Thus, STP implements a frequency-dependent phase preference rather than a simple onset detector. Furthermore, if STP parameters such as release probability $U$ are modified by learning or neuromodulation, phase selectivity shifts accordingly, generating heterogeneity across synapses in which temporal relationships are preferentially learned.
@@ 4.4 Assumptions on Short-Term Facilitation
In subsection 2.2, we neglected short-term facilitation (STF) and modeled short-term plasticity (STP) solely as short-term depression (STD). Experimentally, the facilitation time constant $\tau_{f}$ is typically on the order of tens to a few hundred milliseconds, whereas the recovery time constant from depression $\tau_{d}$ is on the order of several hundred milliseconds up to seconds[zucker2002Shortterm, regehr2012ShortTerm]. Because the associative STP we aim to model is induced on a timescale of at least seconds [kasai2023Mechanical, ucar2021Mechanical], the interaction we primarily need to capture is that between associative plasticity and the slowly recovering depressive component.
@@ 4.5 Limitations and future directions
Several limitations of our current framework warrant discussion and motivate future extensions. First, our analysis employs a perturbative approach, examining small deviations from baseline activity patterns. This limitation prevents us from capturing the full recursive dynamics where recurrent inputs reshape representations, which in turn modify learning rules, creating complex feedback loops. The interaction between synaptic dynamics and neural dynamics likely becomes particularly important when depression timescales approach those of neural integration, potentially leading to rich computational regimes not captured by our linear analysis [hansel2013Shortterm, ocker2015SelfOrganization].
@@
Third, we model spiking as an inhomogeneous Poisson process without refractoriness or reset mechanism, which yields a tractable link between filtered synaptic input and instantaneous firing rate. This approximation is most appropriate in regimes where spiking is highly irregular and the relevant stimulus/synaptic timescales are slow compared with the refractory timescale, such that spike-history effects contribute only weak corrections to rate-based statistics and information measures. Outside this regime, however, real neurons exhibit absolute/relative refractoriness and other spike-history dependencies that induce negative serial correlations and can interact with short-term depression in complex ways [schwalger2017Theory]. These history-dependent effects may therefore modify both Fisher information and the optimal plasticity updates predicted by our framework. A natural future direction is to extend the model to a more general conditional-intensity (escape-rate/GLM) formulation with an explicit spike-history term (or renewal/dead-time dynamics), and re-derive the Fisher information gradients under this enriched point-process description.
@@
Despite these limitations, our framework is readily extensible to address many of these challenges. The mathematical formalism developed here can accommodate different models of presynaptic dynamics, including short-term facilitation-dominant regimes, which may yield qualitatively different learning rules and representation strategies. Similarly, extending the analysis to alternative objective functions – such as mutual information [toyoizumi2005Generalized, toyoizumi2005Spiketiming, pfister2006Optimal], sparse coding [olshausen1996Emergence], or predictive coding [rao1999Predictive] – would be promising within our framework.
@@ 5 Conclusion
Appendix A Exact gradient formula
@@
A.1 Point-process likelihood and the score-function identity
@@
$J(Z)=\left\langle\mathcal{J}[X,Z]\right\rangle_{X(T)}=\int\mathcal{J}[X,Z]\,P(X(T)\mid Z)\,dX,$ (A.2)
@@
$\frac{\partial J}{\partial Z}=\left\langle\frac{\partial\mathcal{J}[X,Z]}{\partial Z}+\mathcal{J}[X,Z]\,\frac{\partial}{\partial Z}\log P(X(T)\mid Z)\right\rangle_{X(T)}.$ (A.3)
Applying (A.3) to $Z=Z_{ij}$ yields (8).
Differentiating (A.1) gives
$\frac{\partial}{\partial Z}\log P(X(T)\mid Z)=\sum_{k=1}^{N}\int_{0}^{T}\left[dN_{k}(t)-\rho_{k}(t)\,dt\right]\,\frac{\partial}{\partial Z}\log\rho_{k}(t).$ (A.4)
@@
$\frac{\partial}{\partial Z}\log\rho_{k}(t)=\frac{g_{k}^{\prime}(t)}{g_{k}(t)}\,\frac{\partial u_{k}(t)}{\partial Z},\qquad g_{k}^{\prime}(t):=\left.\frac{dg}{du}\right|_{u=u_{k}(t)}.$ (A.5)
For the synapse-specific parameter $Z_{ij}\in\{w_{ij}^{0},U_{ij}\}$, the membrane potential $u_{k}(t)$ depends on $Z_{ij}$ directly only when $k=i$. Therefore (A.4)–(A.5) reduce to
$\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)=\int_{0}^{T}\left[dN_{i}(t)-\rho_{i}(t)\,dt\right]\,\frac{g_{i}^{\prime}(t)}{g_{i}(t)}\,\underbrace{\frac{\partial u_{i}(t)}{\partial Z_{ij}}}_{=:~e_{ij}^{Z}(t)},$ (A.6)
@@
A.2 Pathwise derivative of the Fisher-information functional
@@
$A_{k}(t):=h_{k}^{\prime}(t,\theta)\frac{g_{k}^{\prime}(t)}{g_{k}(t)},\qquad\rho_{k}(t)=g_{k}(t):=g(u_{k}(t)).$ (A.7)
@@
$\frac{\partial\mathcal{J}[X]}{\partial Z_{ij}}$ $=\int_{0}^{T}dt\;\frac{\partial}{\partial u_{i}}\left(h_{i}^{\prime}(t,\theta)^{2}\,\frac{g_{i}^{\prime}(t)^{2}}{g_{i}(t)}\right)\,\frac{\partial u_{i}(t)}{\partial Z_{ij}}.$ (A.9)
@@
$\frac{\partial}{\partial u}\left(\frac{g^{\prime}(u)^{2}}{g(u)}\right)=\frac{g^{\prime}(u)^{2}}{g(u)}\left(\frac{2g^{\prime\prime}(u)}{g^{\prime}(u)}-\frac{g^{\prime}(u)}{g(u)}\right).$ (A.10)
@@
$\frac{\partial\mathcal{J}[X]}{\partial Z_{ij}}$ $=\int_{0}^{T}dt\;\rho_{i}(t)\,\eta_{i}(t)\,\underbrace{\frac{\partial u_{i}(t)}{\partial Z_{ij}}}_{=:~e_{ij}^{Z}(t)}.$ (A.11)
@@
A.3 A note on unbiased estimation and variance reduction
Combining (8), (A.6), and the pathwise term yields an unbiased estimator from simulated trajectories:
$\begin{split}\widehat{\frac{\partial J}{\partial Z_{ij}}}&=\frac{1}{M}\sum_{m=1}^{M}\Bigg[\int_{0}^{T}dt\;\rho_{i}^{(m)}(t)\eta_{i}^{(m)}(t)e_{ij}^{Z,(m)}(t)\\ &+\mathcal{J}[X^{(m)}]\int_{0}^{T}\left(dN_{i}^{(m)}(t)-\rho_{i}^{(m)}(t)\,dt\right)\frac{g_{i}^{\prime,(m)}(t)}{g_{i}^{(m)}(t)}e_{ij}^{Z,(m)}(t)\Bigg].\end{split}$ (A.12)

### v2

@@ 4 Discussion
We derived Fisher-information-maximizing learning rules for synapses with Tsodyks–Markram dynamics, allowing both baseline synaptic strength and release probability to be optimized. The resulting learning rules exhibit a temporal bias that preferentially strengthens anti-causal connections, where presynaptic neurons lag behind postsynaptic neurons. In recurrent circuits, this bias reshapes stimulus-evoked activity into ramping representations that emphasize stimulus offsets and supports reverse replay sequences during spontaneous activity after stimulus removal. We also showed that the strength of temporal asymmetry is systematically modulated by global constraints on release probability, suggesting a resource-based mechanism that could tune the balance between forward and reverse replay across behavioral states.
@@ 4.1 Pre- versus postsynaptic plasticity: Computational division of labor
The segregation of synaptic plasticity mechanisms between pre- and postsynaptic sites has long served as a guiding principle for understanding synaptic computation. Classical long-term potentiation is mediated primarily by postsynaptic mechanisms, notably changes in AMPA receptor trafficking and expression (?). By contrast, short-term plasticity has typically been modeled as a predominantly presynaptic phenomenon governed by vesicle dynamics and calcium-dependent modulation of release probability (?, ?). This division of labor has been interpreted as separating associative learning (postsynaptic, detecting pre-post correlations) from adaptive filtering (presynaptic, responding only to presynaptic history).
Recent theoretical work has argued that this pre–post division reflects a deeper computational separation rather than a purely biophysical contingency. Postsynaptic mechanisms, operating on longer timescales, are well suited for stable memory storage and the formation of neural assemblies that encode persistent representations (?, ?). Presynaptic mechanisms, with their rapid dynamics, can implement adaptive filters that extract temporal features from spike trains, enhance signal-to-noise ratios, and maintain homeostatic balance in network activity (?, ?).
The discovery of associative short-term plasticity by Ucar et al. (?) challenges this traditional dichotomy. Their experiments demonstrate that presynaptic release probability can be modulated via postsynaptic spine enlargement, in a manner that requires coincident pre- and postsynaptic activity. This mechanism therefore combines the associative detection capabilities conventionally attributed to postsynaptic plasticity with the rapid, history-dependent filtering characteristic of presynaptic dynamics. Our theoretical framework indicates that such a hybrid rule is not simply “LTP made faster”, but gives rise to distinct forms of temporal selectivity.
@@ 4.2 Reverse replay and state-dependent plasticity
This mechanism aligns with and extends the model proposed by (?), which demonstrated that symmetric STDP combined with short-term depression can generate backward connectivity supporting reverse replay. In our framework, such connectivity patterns arise from an information-theoretic optimization principle: a backward bias is not an incidental artifact of a particular plasticity rule, but an expected outcome when information transmission is optimized under appropriate constraints.
@@
We interpret this “release-probability budget” as an effective parameter that can be influenced by physiological mechanisms such as neuromodulation. Neuromodulators vary strongly across behavioral states and sleep stages and modulate presynaptic properties, including short-term dynamics (?, ?). Recent work further underscores that presynaptic resources are regulated at the behavioral-state level: Wu et al. (?) showed across multiple circuits in Drosophila and mice that presynaptic release probability decreases with prolonged wakefulness and recovers after sleep, and that bidirectional manipulation of release probability bidirectionally alters sleep need. Complementary studies in Drosophila demonstrate that presynaptic active zone composition, which closely tracks release probability, is remodeled as a function of sleep pressure (?). Although these findings do not address hippocampal replay directly, they establish that presynaptic resource constraints can indeed be tuned in a state-dependent manner, consistent with the key premise of our model.
@@
This picture dovetails with empirical observations that reverse replay is especially prominent during quiet wakefulness following spatial experience and reward delivery, whereas forward replay is more prevalent during sleep sharp-wave ripples (?, ?, ?, ?, ?). In our framework, these regimes correspond to different effective presynaptic budgets acting on top of a fixed forward backbone: more generous budgets promote the emergence of reverse replay that may facilitate rapid behavioral adaptation and reward credit assignment (?), whereas more stringent budgets yield connectivity that expresses predominantly forward replay and may better support stabilization and generalization of memories during sleep (?, ?, ?, ?). This proposal links state-dependent neuromodulation of presynaptic function to a quantitative, information-theoretic control knob for the balance between forward and reverse replay, and thereby frames reverse sequences not as an oddity of hippocampal dynamics, but as a predicted feature of efficient temporal coding under varying resource constraints.
@@ 4.3 Frequency-Dependent Phase Selectivity in Pre- and Postsynaptic Channels
Our linear-response analysis of $C^{Z}(t)$ (Figure 2; Appendix 5.4.2) confirms that onset selectivity is robust, but the phase of maximal plasticity shifts with input frequency and the neuron–synapse operating point. Thus, STP implements a frequency-dependent phase preference rather than a simple onset detector. Furthermore, if STP parameters such as release probability $U$ are modified by learning or neuromodulation, phase selectivity shifts accordingly, generating heterogeneity across synapses in which temporal relationships are preferentially learned.
@@ 4.4 Assumptions on Short-Term Facilitation
In subsection 2.2, we neglected short-term facilitation (STF) and modeled short-term plasticity (STP) solely as short-term depression (STD). Experimentally, the facilitation time constant $\tau_{f}$ is typically on the order of tens to a few hundred milliseconds, whereas the recovery time constant from depression $\tau_{d}$ is on the order of several hundred milliseconds up to seconds (?, ?). Because the associative STP we aim to model is induced on a timescale of at least seconds (?, ?), the interaction we primarily need to capture is that between associative plasticity and the slowly recovering depressive component.
@@ 4.5 Limitations and future directions
Several limitations of our current framework warrant discussion and motivate future extensions.
@@
Second, we model spiking as an inhomogeneous Poisson process without refractoriness or reset mechanism, which yields a tractable link between filtered synaptic input and instantaneous firing rate. This approximation is most appropriate in regimes where spiking is highly irregular and the relevant stimulus/synaptic timescales are slow compared with the refractory timescale, such that spike-history effects contribute only weak corrections to rate-based statistics and information measures. Outside this regime, however, real neurons exhibit absolute/relative refractoriness and other spike-history dependencies that induce negative serial correlations and can interact with short-term depression in complex ways (?). These history-dependent effects may therefore modify both Fisher information and the optimal plasticity updates predicted by our framework. A natural future direction is to extend the model to a more general conditional-intensity (escape-rate/GLM) formulation with an explicit spike-history term (or renewal/dead-time dynamics), and re-derive the Fisher information gradients under this enriched point-process description.
@@
Despite these limitations, our framework is readily extensible to address many of these challenges. The mathematical formalism developed here can accommodate different models of presynaptic dynamics, including short-term facilitation-dominant regimes, which may yield qualitatively different learning rules and representation strategies. Similarly, extending the analysis to alternative objective functions – such as mutual information (?, ?, ?), sparse coding (?), or predictive coding (?) – would be promising within our framework.
@@ 5 Conclusion
Figure 7: Learning-induced changes in stimulus-evoked activity. Changes in neural representations induced by Fisher-information gradient learning. A. Associative STP. B. Non-associative STP: synapses obey Tsodyks–Markram dynamics, but the release probability $U(\zeta)$ is fixed and only $w_{0}(\zeta)$ is learned. C. Static synapses (no STP dynamics; synaptic weights are fixed during simulation). In B, we fix $U(\zeta)=0.15$. Other parameters: $A=3.0$, $\theta_{c}=0$, $\tau_{d}=0.5~\mathrm{s}$, $\beta=2.0$, $g_{c}=10.0$, $u_{c}=1.0$, $\tau_{s}=0.01~\mathrm{s}$.
Figure 8: Cross-correlograms during evoked and spontaneous activity. Pairwise spike-train cross-correlograms after Fisher-information optimization. A–C. During stimulus presentation. D–F. During spontaneous activity. A,D. Associative STP. B,E. Non-associative STP. C,F. Static synapses. same params as previous figure.
Figure 9: Emergence of reverse replay. Spontaneous activity after learning to optimize Fisher information. Reverse replay is observed in associative STP (A) and non-associative STP (B), whereas static synapses show no reliable replay (C). Under associative STP, replay is more temporally compressed (faster propagation) and exhibits sharper spatial contrast than under non-associative STP. A. Associative STP. B. Non-associative STP. C. Static synapses. same params as previous figure.
Acknowledgments
The author gratefully acknowledges Kensuke Yoshida and all members of the Toyoizumi Lab for insightful advice and stimulating discussions. Heartfelt gratitude to my wife and newborn child for their unwavering support and inspiration.
Funding:
G. S. was funded by JSPS KAKENHI Grant Number 23KJ0666. T. T. was funded by JST CREST JPMJCR23N2, RIKEN Center for Brain Science, and RIKEN TRIP initiate (RIKEN Quantum).
Author contributions:
G.S. performed the research, analyzed the data, and wrote the manuscript. T.T. supervised the project and provided critical feedback. Both authors approved the final version of the manuscript.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials.
Supplementary materials
Materials and Methods
Figs. S1 to S8
Supplementary Materials for
Genki Shimizu, Taro Toyoizumi
Corresponding author. Email: genki.shimizu@a.riken.jp
This PDF file includes:
Materials and Methods
Figs. S1 to S8
Materials and Methods
5.1 Exact gradient formula
@@
5.1.1 Point-process likelihood and the score-function identity
@@
$J(Z)=\left\langle\mathcal{J}[X,Z]\right\rangle_{X(T)}=\int\mathcal{J}[X,Z]\,P(X(T)\mid Z)\,dX,$ (S2)
@@
$\frac{\partial J}{\partial Z}=\left\langle\frac{\partial\mathcal{J}[X,Z]}{\partial Z}+\mathcal{J}[X,Z]\,\frac{\partial}{\partial Z}\log P(X(T)\mid Z)\right\rangle_{X(T)}.$ (S3)
Applying (S3) to $Z=Z_{ij}$ yields (8).
Differentiating (S1) gives
$\frac{\partial}{\partial Z}\log P(X(T)\mid Z)\\ =\sum_{k=1}^{N}\int_{0}^{T}\left[dN_{k}(t)-\rho_{k}(t)\,dt\right]\,\frac{\partial}{\partial Z}\log\rho_{k}(t).$ (S4)
@@
$\frac{\partial}{\partial Z}\log\rho_{k}(t)=\frac{g_{k}^{\prime}(t)}{g_{k}(t)}\,\frac{\partial u_{k}(t)}{\partial Z},\qquad g_{k}^{\prime}(t):=\left.\frac{dg}{du}\right|_{u=u_{k}(t)}.$ (S5)
For the synapse-specific parameter $Z_{ij}\in\{w_{ij}^{0},U_{ij}\}$, the membrane potential $u_{k}(t)$ depends on $Z_{ij}$ directly only when $k=i$. Therefore (S4)–(S5) reduce to
$\frac{\partial}{\partial Z_{ij}}\log P(X(T)\mid Z)\\ =\int_{0}^{T}\left[dN_{i}(t)-\rho_{i}(t)\,dt\right]\,\frac{g_{i}^{\prime}(t)}{g_{i}(t)}\,\underbrace{\frac{\partial u_{i}(t)}{\partial Z_{ij}}}_{=:~e_{ij}^{Z}(t)},$ (S6)
@@
5.1.2 Pathwise derivative of the Fisher-information functional
@@
$A_{k}(t):=h_{k}^{\prime}(t,\theta)\frac{g_{k}^{\prime}(t)}{g_{k}(t)},\qquad\rho_{k}(t)=g_{k}(t):=g(u_{k}(t)).$ (S7)
@@
$\frac{\partial\mathcal{J}[X]}{\partial Z_{ij}}$ $=\int_{0}^{T}dt\;\frac{\partial}{\partial u_{i}}\left(h_{i}^{\prime}(t,\theta)^{2}\,\frac{g_{i}^{\prime}(t)^{2}}{g_{i}(t)}\right)\,\frac{\partial u_{i}(t)}{\partial Z_{ij}}.$ (S9)
@@
$\frac{\partial}{\partial u}\left(\frac{g^{\prime}(u)^{2}}{g(u)}\right)=\frac{g^{\prime}(u)^{2}}{g(u)}\left(\frac{2g^{\prime\prime}(u)}{g^{\prime}(u)}-\frac{g^{\prime}(u)}{g(u)}\right).$ (S10)
@@
$\frac{\partial\mathcal{J}[X]}{\partial Z_{ij}}$ $=\int_{0}^{T}dt\;\rho_{i}(t)\,\eta_{i}(t)\,\underbrace{\frac{\partial u_{i}(t)}{\partial Z_{ij}}}_{=:~e_{ij}^{Z}(t)}.$ (S11)
@@
5.1.3 A note on unbiased estimation and variance reduction
Combining (8), (S6), and the pathwise term yields an unbiased estimator from simulated trajectories:
$\begin{split}\widehat{\frac{\partial J}{\partial Z_{ij}}}&=\frac{1}{M}\sum_{m=1}^{M}\Bigg[\int_{0}^{T}dt\;\rho_{i}^{(m)}(t)\eta_{i}^{(m)}(t)e_{ij}^{Z,(m)}(t)\\ &+\mathcal{J}[X^{(m)}]\int_{0}^{T}\left(dN_{i}^{(m)}(t)-\rho_{i}^{(m)}(t)\,dt\right)\frac{g_{i}^{\prime,(m)}(t)}{g_{i}^{(m)}(t)}e_{ij}^{Z,(m)}(t)\Bigg].\end{split}$ (S12)